OpenStudy (stormoshawty):

Write an equation in slope-intercept, point-slope, or standard form for the line with the given information. Explain why you chose the form you did. A. passes through (-1, 4) and (-5, 2) B. slope 2, y-intercept -4 C. has an x-intercept of 6 and a y-intercept of 3 D. passes through (1, 2) with a slope of - 5/3

OpenStudy (stormoshawty):

@shamil98

OpenStudy (anonymous):

do you know the forms its asking about?

OpenStudy (stormoshawty):

Yes.

OpenStudy (anonymous):

so write your forms of equations down and look for variables that you have values for, for each equation.

OpenStudy (stormoshawty):

Omg .-.

OpenStudy (stormoshawty):

That doesn't help me at all, that's kinda WHY i posted it on here.

OpenStudy (stormoshawty):

@Luigi0210 @Squirrels

OpenStudy (anonymous):

so you basically want the answer? if you have 2 points you dont want to use slope intercept or standard form do you

OpenStudy (stormoshawty):

Yeah, that's the only reason anyone posts on here, is for answers.

OpenStudy (anonymous):

well that doesn't do you any good to just get the answer. math is all about the process to get the answer. how will you do this on an exam if you just get the answer

OpenStudy (stormoshawty):

I already know how to do all of this stuff, i'm just in a hurry to get it done, so i know how, i just want it done now.

OpenStudy (nincompoop):

A. passes through (-1, 4) and (-5, 2) y = mx + b m=? \[m=\frac{ \Delta y }{ \Delta x }=\frac{ y_2-y_1 }{ x_2-x_1 }\] \[m=\frac{ \Delta y }{ \Delta x }=\frac{ y_2-y_1 }{ x_2-x_1 }=\frac{ 2-4 }{-5-1 }=\frac{ 1 }{ 3 }\] \[y-4=\frac{ 1 }{ 3 }(x--1)\] \[y=\frac{ 1 }{ 3 }x+3\frac{ 2 }{ 3 }\]

OpenStudy (nincompoop):

B. slope 2, y-intercept -4 m = 2 ; (0, -4) y= 2x -4

OpenStudy (nincompoop):

C. has an x-intercept of 6 and a y-intercept of 3 x-intercept (6, 0) y-intercept (0, 3) y = mx + b \[m=\frac{ \Delta y }{ \Delta x }=\frac{ y_2-y_1 }{ x_2-x_1 }=\frac{ 3-0 }{ 0-6 }=-\frac{ 1 }{ 2 }\] \[y = \frac{ 1 }{ 2 }x+3\]

OpenStudy (stormoshawty):

Is B slope-intercept or standard?

OpenStudy (nikato):

Slope intercept. U already have the intercept and the slope

OpenStudy (nincompoop):

D. passes through (1, 2) with a slope of - 5/3 \[y-y_1 = m(x-x_1)\] \[y-2=-\frac{ 5 }{ 3 }x-\frac{ 5 }{ 3 }\] \[y = -\frac{ 5 }{ 3 }x-\frac{ 5 }{ 3 }+\frac{ 6 }{ 3 }=-\frac{ 5}{ 3 }x+\frac{ 1 }{ 3 }\] y = 5x/3 + 1/3

OpenStudy (stormoshawty):

What form did you use for D?

OpenStudy (stormoshawty):

Nevermind.

OpenStudy (stormoshawty):

So a-c is slope-intercept?

OpenStudy (nincompoop):

they're all in slope-intercept form you said that you know all of them already

OpenStudy (nincompoop):

it would have taken you the same amount of time to solve all of the questions as I have, so I am not really sure why you're looking for answers if you already know this stuff

OpenStudy (stormoshawty):

D isn't slope-intercept. D is point-slope.

OpenStudy (nincompoop):

point-slope if you are given one coordinate point and a slope slope-intercept form is one of the "forms" of an equation of a line the general form is ax+by=c

OpenStudy (nikato):

Yea, D is point slope

OpenStudy (nincompoop):

an explanation is not required @nikato she assured us that she knew the concept already

OpenStudy (nikato):

I know @nincompoop. Maybe she just wanted to check her answers. Btw, no offense, but ur profile pic really creeps me out

OpenStudy (nincompoop):

don't be judgmental

OpenStudy (nikato):

I'm sorry, but it just really creeps me out.

OpenStudy (nincompoop):

Next time, I will make sure to point out that this website is not where you go to look for an answer. this is a website to learn those things you couldn't grasp in your lessons. I usually go over concepts using different sets of problems instead of going by the homework or assignment. IF for some reason you ask more questions about line problems, it will mean that you lied to us that you knew about this stuff already.